# Multigrade Neural Network Approximation

### v1

@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable empirical success in areas such as computer vision, natural language processing, and scientific computing, owing to their strong expressive power for modeling complex input-output relationships. From an approximation-theoretic perspective, this expressive capacity is now relatively well understood: sufficiently deep and wide networks can approximate broad classes of functions with high accuracy [38, 32, 48]. Despite these advances, the effective training of deep neural networks with many layers is still not fully understood. Standard end-to-end training leads to highly non-convex and often ill-conditioned optimization landscapes, making performance sensitive to initialization, learning rates, and architectural choices. These challenges are further exacerbated by vanishing or exploding gradients, spectral bias toward low-frequency features [27, 47], and short-term oscillations near the edge of stability [1, 6]. In contrast, the optimization landscape of shallow architectures is fundamentally different; in particular, training one-hidden-layer $\mathtt{ReLU}$ networks can be reformulated as convex optimization problems, yielding global optimality and strong theoretical guarantees for both training and generalization [26, 11].
@@ 1.1 Multigrade deep learning
Figure 1: Overview of the MGDL framework and its grade-wise training procedure. Training proceeds grade by grade: at each grade, only the newly added block and the output map are optimized, while all previously learned blocks are frozen and serve as adaptive basis components. The network $\Phi_{m}$ approximates the target function via a recursive residual refinement process.
@@
Thus, each grade solves a well-defined subproblem: $T_{m}$ extracts new features based on all preceding transformations, and $g_{m}$ fits the current residual using these features. A detailed algorithmic formulation of this grade-wise optimization procedure is provided in Section 2.
@@ 1.2 Main theorem
    For any preassigned $p\in[1,\infty)$, $\|R_{k}\|_{L^{p}([0,1]^{d})}\searrow 0$ strictly as $k\to\infty$.
@@ 1.3 Understanding MGDL from dictionary-based approximation
Sparse approximation adopts a more flexible, but technically more involved, approach. Instead of relying on a basis, it employs a dictionary, typically a redundant system, such as multiresolution-analysis-based wavelet tight frames in $L^{2}(\mathbb{R}^{d})$ [29]. Given a dictionary $\mathcal{D}=\{{\bm{e}}_{i}\}_{i\in\mathcal{I}}$, the goal is to approximate a target function using only a small number of dictionary elements. A standard greedy strategy proceeds by iteratively reducing the residual. Starting from $R_{0}=f$, one selects at each iteration an atom–coefficient pair by solving
@@
More precisely, MGDL constructs the compositional dictionary
[…]
For notational convenience, we denote by $\mathcal{D}_{\mathrm{MGDL}}(k)$ the subset of $\mathcal{D}_{\mathrm{MGDL}}$ consisting of compositions of the form $T_{k}\circ\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu}\circ T_{1}$.
Starting from $R_{0}=f$, MGDL proceeds iteratively by selecting, at each grade, an atom together with an output map $g\in\mathcal{G}$ through the optimization problem
$(g_{k+1},{\bm{e}}_{k+1})\;\in\;\operatorname*{arg\,min}_{g\in\mathcal{G},\,{\bm{e}}\in\mathcal{D}_{\mathrm{MGDL}}(k)}\|R_{k}-g\circ{\bm{e}}\|.$
@@ 3.1 Proof strategy
Let $f\in C([0,1]^{d})$ be given. There exist $n\in{\mathbb{N}}^{+}$ and affine maps $\mathcal{A}_{i},\,\mathcal{A}_{i}^{\mathrm{out}}\in{\mathsf{Aff}}_{\leq 5d}$ for $i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},n$ such that, defining $f_{1}\coloneqq f$ and
@@ 3.2 Main argument
Apply Theorem 3.1 to $f_{1}\coloneqq f\in C([0,1]^{d})$. There exist $n_{1}\in{\mathbb{N}}^{+}$ and affine maps $\mathcal{A}_{i},\mathcal{A}_{i}^{\mathrm{out}}\in{\mathsf{Aff}}_{\leq 5d}$ for $i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},n_{1}$ such that, with
@@
Next, apply Theorem 3.1 again to $f_{n_{1}+1}\in C([0,1]^{d})$. There exist $n_{2}\in{\mathbb{N}}^{+}$ and affine maps $\mathcal{A}_{k},\mathcal{A}_{k}^{\mathrm{out}}\in{\mathsf{Aff}}_{\leq 5d}$ for $k=n_{1}+1,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},n_{1}+n_{2}$ such that, with
@@
Consequently, all properties established for $\{f_{k}\}_{k=1}^{\infty}$ immediately carry over to the residual sequence $\{R_{k}\}_{k=1}^{\infty}$. In particular, we have

### v2

@@ 1 Introduction
Deep neural networks have achieved remarkable empirical success in areas such as computer vision, natural language processing, and scientific computing, owing to their strong expressive power for modeling complex input-output relationships. From an approximation-theoretic perspective, this expressive capacity is now relatively well understood: sufficiently deep and wide networks can approximate broad classes of functions with high accuracy [38, 32, 48]. Despite these advances, the effective training of deep neural networks with many layers is still not fully understood. Standard end-to-end training leads to highly non-convex and often ill-conditioned optimization landscapes, making performance sensitive to initialization, learning rates, and architectural choices. These challenges are further exacerbated by vanishing or exploding gradients, spectral bias toward low-frequency features [27, 47], and short-term oscillations near the edge of stability [1, 6]. In contrast, the optimization landscape of shallow architectures is fundamentally different; in particular, training one-hidden-layer $\mathtt{ReLU}$ networks can be reformulated as convex optimization problems, yielding global optimality and strong theoretical guarantees for both training and generalization [26, 11].
@@ 1.1 Multigrade deep learning
Figure 1: Overview of the framework and its grade-wise training procedure. Training proceeds grade by grade: at each grade, only the newly added block and the output map are optimized, while all previously learned blocks are frozen and serve as adaptive basis components. The network $\Phi_{m}$ approximates the target function via a recursive residual refinement process.
@@
Thus, each grade solves a well-defined subproblem: $T_{m}$ extracts new features based on all preceding transformations, and $g_{m}$ fits the current residual using these features. MGDL adopts a stage-wise learning strategy, where each grade searches and optimizes only within a relatively low-complexity parameter (hypothesis) space; see Figure 2 for an illustration of this grade-wise training process. This design keeps each optimization step simple and interpretable, while raising a central theoretical issue about its overall approximation capability. The main goal of this work is to show that, under appropriate assumptions, MGDL’s grade-wise training produces an approximation error that decreases monotonically and converges to zero as the number of grades tends to infinity. A detailed algorithmic formulation of this grade-wise optimization procedure is provided in Section 2.
Figure 2: Overview of the MGDL grade-wise training procedure.
@@ 1.2 Main theorem
    Given any $p\in[1,\infty)$, $\|R_{k}\|_{L^{p}([0,1]^{d})}\searrow 0$ strictly until it reaches zero as $k\to\infty$, i.e., for any $k$, either $R_{k+1}\equiv R_{k}\equiv 0$ on $0,1]^{d}$ or $\|R_{k+1}\|_{L^{p}([0,1]^{d})}<\|R_{k}\|_{L^{p}([0,1]^{d})}$.
@@ 1.3 Understanding MGDL from dictionary-based approximation
Sparse approximation adopts a more flexible, but technically more involved, approach. Instead of relying on a basis, it employs a dictionary, typically a redundant system, such as multiresolution-analysis-based wavelet tight frames in $L^{2}(\mathbb{R}^{d})$ [29]. Given a dictionary $\mathcal{D}=\{{\bm{e}}_{i}\}_{i\in\mathcal{I}}$, the goal is to approximate a target function using only a small number of dictionary elements. A standard greedy strategy proceeds by iteratively reducing the residual. Starting from $R_{0}=f$, one selects at each iteration an atom-coefficient pair by solving
@@
Starting from $R_{0}=f$ and the initial dictionary
$\mathcal{D}_{\mathrm{MGDL}}(1)=\{T_{1}:T_{1}\text{ is a network block}\},$
MGDL proceeds iteratively by selecting, at each grade, an atom together with an output map $g\in\mathcal{G}$ through the optimization problem
$(g_{k+1},{\bm{e}}_{k+1})\;\in\;\operatorname*{arg\,min}_{g\in\mathcal{G},\,{\bm{e}}\in\mathcal{D}_{\mathrm{MGDL}}(k+1)}\|R_{k}-g\circ{\bm{e}}\|,$
where the new dictionary $\mathcal{D}_{\mathrm{MGDL}}(k+1)$ is constructed from the atom $\mathbf{e}_{k}$ learned at the previous grade, i.e.,
$\mathcal{D}_{\mathrm{MGDL}}(k+1)=\bigl\{T_{k+1}\circ\mathbf{e}_{k}:T_{k+1}\text{ is a network block}\bigr\}.$
@@ 3.1 Proof strategy
Let $f\in C([0,1]^{d})$ be a nonzero function. There exist $n\in{\mathbb{N}}^{+}$ and affine maps $\mathcal{A}_{i},\,\mathcal{A}_{i}^{\mathrm{out}}\in{\mathsf{Aff}}_{\leq 5d}$ for $i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},n$ such that, defining $f_{1}\coloneqq f$ and
@@ 3.2 Main argument
We assume $f\not\equiv 0$, as the case $f\equiv 0$ is trivial. Apply Theorem 3.1 to $f_{1}\coloneqq f\in C([0,1]^{d})$. There exist $n_{1}\in{\mathbb{N}}^{+}$ and affine maps $\mathcal{A}_{i},\mathcal{A}_{i}^{\mathrm{out}}\in{\mathsf{Aff}}_{\leq 5d}$ for $i=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},n_{1}$ such that, with
@@
Next, we assume $f_{n_{1}+1}\not\equiv 0$, since the case $f_{n_{1}+1}\equiv 0$ is trivial. Applying Theorem 3.1 again to $f_{n_{1}+1}\in C([0,1]^{d})$, there exist $n_{2}\in{\mathbb{N}}^{+}$ and affine maps $\mathcal{A}_{k},\mathcal{A}_{k}^{\mathrm{out}}\in{\mathsf{Aff}}_{\leq 5d}$ for $k=n_{1}+1,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},n_{1}+n_{2}$ such that, with
@@
We remark that in the above construction, if $R_{k_{0}}=f_{k_{0}+1}\equiv 0$ for some $k_{0}$, then it suffices to define $\Phi_{k+1}\equiv 0$ for all $k\geq k_{0}$, which implies that $R_{k}\equiv 0$ for all $k\geq k_{0}$.
Clearly, all properties established for $\{f_{k}\}_{k=1}^{\infty}$ immediately carry over to the residual sequence $\{R_{k}\}_{k=1}^{\infty}$. In particular, we have
@@ 4.3 Main argument
We remark that $n_{f,\varepsilon}$ and $n_{-f,\varepsilon}$ may be zero. For instance, $n_{f,\varepsilon}=0$ if $f(\bm{x})\leq 0$ for all $\bm{x}\in[0,1]^{d}$.